# INN-PAR: Invertible Neural Network for PPG to ABP Reconstruction

### v1

@@ Abstract
Non-invasive and continuous blood pressure (BP) monitoring is essential for the early prevention of many cardiovascular diseases. Estimating arterial blood pressure (ABP) from photoplethysmography (PPG) has emerged as a promising solution. However, existing deep learning approaches for PPG-to-ABP reconstruction (PAR) encounter certain information loss, impacting the precision of the reconstructed signal. To overcome this limitation, we introduce an invertible neural network for PPG to ABP reconstruction (INN-PAR), which employs a series of invertible blocks to jointly learn the mapping between PPG and its gradient with the ABP signal and its gradient. INN-PAR efficiently captures both forward and inverse mappings simultaneously, thereby preventing information loss. By integrating signal gradients into the learning process, INN-PAR enhances the network’s ability to capture essential high-frequency details, leading to more accurate signal reconstruction. Moreover, we propose a multi-scale convolution module (MSCM) within the invertible block, enabling the model to learn features across multiple scales effectively. We have experimented on two benchmark datasets, which show that INN-PAR significantly outperforms the state-of-the-art methods in both waveform reconstruction and BP measurement accuracy.
@@ I Introduction
Though INNs have primarily been used for generative tasks in image [8, 10, 11] and signal processing [12, 13, 14], several studies have also applied INNs to image reconstruction tasks [15, 16, 17, 18]. In a similar manner, INN may be advantageous for signal-to-signal reconstruction tasks, as they simultaneously learn both the forward and inverse mappings, which helps prevent information loss. Additionally, incorporating the signal gradient into the learning process, which focuses on the rate of change rather than just the signal values, enhances the ability to capture essential high-frequency details [19, 20], thereby ensuring a more accurate reconstruction of the signal’s shape.
@@ III-A Experimental Setup
For the experiments, we utilize two publicly available datasets: Sensors [2] and BCG [21]. We use the Sensors dataset for training and testing INN-PAR. Moreover, we also tested the performance of the BCG dataset to check the model’s generalization ability without any finetuning. We follow [1] for data pre-processing and splitting the dataset into training, validation, and test sets. We are not using any subject calibration and PPG scaling. Also, we do not allow subject information leakage among the training, validation, and test sets. The signals are segmented into $5s$ chunks without overlapping. For the Sensors dataset, we use $6658$ signal chunks for training, $2383$ for validation, and $2061$ for testing. For the BCG dataset, we use $671$ signal chunks for testing. We train INN-PAR by minimizing the loss function in Eqn. 5 using Adam optimizer ($\beta_{1}=0.9,\>\beta_{2}=0.999$) for $500$ epochs with a batch size of $128$. The learning rate is kept constant at $1\times 10^{-4}$. Using the Pytorch framework, we have conducted all the experiments in the NVIDIA A40 GPU. Mean absolute error (MAE) and normalized root mean square error (NRMSE) [23] metrics are used to compare the waveform reconstruction accuracy. For the accuracy of SBP and DBP measurement, we utilize the MAE metric.
@@ III-B Performance Comparison
Figure 3 shows the visual comparison for a signal from the Sensors and BCG datasets. VNet1d does not reconstruct the ABP signal well. UNet1d and INN-PAR show better signal reconstruction capability. However, INN-PAR reconstructs the high-frequency details of the ABP signal with much better accuracy. All these quantitative and qualitative results show the effectiveness of INN-PAR.
@@ IV Conclusion
This work introduces INN-PAR, an invertible network designed for the PPG-to-ABP reconstruction (PAR) task. Unlike other deep learning models, INN-PAR jointly learns the mapping between PPG and its gradient with those of the ABP signal, simultaneously capturing both forward and inverse mappings, thus, preventing information loss. Also, we propose MSCM to capture features at multiple scales. Our ablation experiments justify the design of INN-PAR. Experimental results demonstrate that INN-PAR outperforms the SOTA methods. Future research could investigate the application of INN-PAR to other physiological signal reconstruction tasks.

### v2

@@ Abstract
Non-invasive and continuous blood pressure (BP) monitoring is essential for the early prevention of many cardiovascular diseases. Estimating arterial blood pressure (ABP) from photoplethysmography (PPG) has emerged as a promising solution. However, existing deep learning approaches for PPG-to-ABP reconstruction (PAR) encounter certain information loss, impacting the precision of the reconstructed signal. To overcome this limitation, we introduce an invertible neural network for PPG to ABP reconstruction (INN-PAR), which employs a series of invertible blocks to jointly learn the mapping between PPG and its gradient with the ABP signal and its gradient. INN-PAR efficiently captures both forward and inverse mappings simultaneously, thereby preventing information loss. By integrating signal gradients into the learning process, INN-PAR enhances the network’s ability to capture essential high-frequency details, leading to more accurate signal reconstruction. Moreover, we propose a multi-scale convolution module (MSCM) within the invertible block, enabling the model to learn features across multiple scales effectively. We have experimented on two benchmark datasets, which show that INN-PAR significantly outperforms the state-of-the-art methods in both waveform reconstruction and BP measurement accuracy. Codes can be found at: https://github.com/soumitra1992/INNPAR-PPG2ABP.
@@ I Introduction
Though INNs have primarily been used for generative tasks in image [8, 10, 11] and signal processing [12, 13, 14], several studies have also applied INNs for image reconstruction tasks [15, 16, 17, 18]. In a similar manner, INNs may be advantageous for signal-to-signal reconstruction tasks, as they simultaneously learn both the forward and inverse mappings, which helps prevent information loss. Additionally, incorporating the signal gradient into the learning process, which focuses on the rate of change rather than just the signal values, enhances the ability to capture essential high-frequency details [19, 20], thereby ensuring a more accurate reconstruction of the signal’s shape.
@@ III-A Experimental Setup
For the experiments, we utilize two publicly available datasets: Sensors [2] and BCG [21]. We use the Sensors dataset for training and testing INN-PAR. Moreover, we also test the performance of INN-PAR on the BCG dataset to check the model’s generalization ability, without any finetuning. We follow [1] for data pre-processing and splitting the dataset into training, validation, and test sets. We are not using any subject calibration and PPG scaling. Also, we do not allow subject information leakage among the training, validation, and test sets. The signals are segmented into $5s$ chunks without overlapping. For the Sensors dataset, we use $6658$ signal chunks for training, $2383$ for validation, and $2061$ for testing. For the BCG dataset, we use $671$ signal chunks for testing. We train INN-PAR by minimizing the loss function in Eqn. 5 using Adam optimizer ($\beta_{1}=0.9,\>\beta_{2}=0.999$) for $500$ epochs with a batch size of $128$. The learning rate is kept constant at $1\times 10^{-4}$. Using the Pytorch framework, we have conducted all the experiments in the NVIDIA A40 GPU. Mean absolute error (MAE) and normalized root mean square error (NRMSE) [23] metrics are used to compare the waveform reconstruction accuracy. For the accuracy of SBP and DBP measurement, we utilize the MAE metric.
@@ III-B Performance Comparison
Figure 3 shows the visual comparison of the PAR task for a signal from the Sensors and BCG datasets. VNet1d does not reconstruct the ABP signal well. UNet1d and INN-PAR show better signal reconstruction capability. However, INN-PAR reconstructs the high-frequency details of the ABP signal with much better accuracy. All these quantitative and qualitative results show the effectiveness of INN-PAR.
@@ IV Conclusion
This work introduces INN-PAR, an invertible network designed for the PAR task. Unlike other deep learning models, INN-PAR jointly learns the mapping between PPG and its gradient with those of the ABP signal, simultaneously capturing both forward and inverse mappings, thus, preventing information loss. Also, we propose MSCM to capture features at multiple scales. Our ablation experiments justify the design of INN-PAR. Experimental results demonstrate that INN-PAR outperforms the SOTA methods. Future research could investigate the application of INN-PAR to other physiological signal reconstruction tasks.